### v2

# CoRA: Collaborative Information Perception by Large Language Model’s Weights for Recommendation

Abstract
Involving collaborative information in Large Language Models (LLMs) is a promising technique for adapting LLMs for recommendation. Existing methods achieve this by concatenating collaborative features with text tokens into a unified sequence input and then fine-tuning to align these features with LLM’s input space. Although effective, in this work, we identify two limitations when adapting LLMs to recommendation tasks, which hinder the integration of general knowledge and collaborative information, resulting in sub-optimal recommendation performance. (1) Fine-tuning LLM with recommendation data can undermine its inherent world knowledge and fundamental competencies, which are crucial for interpreting and inferring recommendation text. (2) Incorporating collaborative features into textual prompts disrupts the semantics of the original prompts, preventing LLM from generating appropriate outputs. In this paper, we propose a new paradigm, CoRA (an acronym for Collaborative LoRA), with a collaborative weights generator. Rather than input space alignment, this method aligns collaborative information with LLM’s parameter space, representing them as incremental weights to update LLM’s output. This way, LLM perceives collaborative information without altering its general knowledge and text inference capabilities. Specifically, we employ a collaborative filtering model to extract user and item embeddings, converting them into collaborative weights with low-rank properties through the collaborative weights generator. We then merge the collaborative weights into LLM’s weights, enabling LLM to perceive the collaborative signals and generate personalized recommendations without fine-tuning or extra collaborative tokens in prompts. Extensive experiments confirm that CoRA effectively integrates collaborative information into LLM, enhancing recommendation performance.
Introduction
Large language models (LLMs) have showcased remarkable performance in a wide range of natural language processing tasks and demonstrated excellent generalization capabilities, offering a promising solution to real-world problems. To leverage the extensive competencies of LLMs, an increasing number of studies are investigating ways to frame recommendation problems in natural language, allowing LLMs to tackle these queries (LLMRec) (Li et al. 2023b; Zhang et al. 2024a; Dai et al. 2023a). Collaborative information, which delineates the co-occurrence patterns in user-item interactions (Zhang et al. 2024b), is critical for the success of conventional collaborative filtering methods. Regardless, in LLM-based recommendation approaches, collaborative information cannot be directly interpreted as text. LLMs’ ineptitude in perceiving this information hampers their performance compared to traditional recommendation models. Consequently, enabling LLMs to comprehend and utilize collaborative information presents a significant challenge.
Figure 1: The performance of Vicuna-7B before and after fine-tuning on Amazon-Book using the prompt in TALLRec. The EM/ROUGE-L scores of generated answers on datasets represent various general and recommendation abilities.
Some approaches are beginning to focus on this issue (Li et al. 2023c; Li, Zhang, and Chen 2022). For instance, Prompt4NR (Zhang and wei Wang 2023) makes the first trial of prompt learning paradigm to convert the News Recommendation task as a language prediction task, where user and item data are wrapped into textual prompts. CoLLM (Zhang et al. 2023) captures collaborative information through an external traditional model and maps it to the input space of LLM, forming user and item embeddings as unique tokens for LLM usage. BinLLM (Zhang et al. 2024b) converts collaborative embeddings into binary value sequences that LLMs can operate on directly, facilitating LLMs’ direct usage of collaborative information in text-like format. LlaRA (Liao et al. 2023) uses a hybrid prompting method that integrates ID embeddings with textual features.
Despite effectiveness, these techniques fail to preserve the LLM’s inherent comprehension and inference for universal and recommendation information when adapting LLMs to recommendation. Our experiments reveal that LLM’s capabilities have been weakened from two perspectives, leading to sub-optimal recommendation performance.
(1) Researchers have discovered that supervised fine-tuning on instructions they have never seen will encourage them to produce hallucinations (Li et al. 2024). We evaluate LLM’s interpreting and inferring capabilities for general and recommendation textual descriptions. For general knowledge, we use Vicuna-7B(Chiang et al. 2023) to generate answers on WikiFact (Goodrich et al. 2019), SocialIQA (Sap et al. 2019), and XSum (Narayan, Cohen, and Lapata 2018) datasets. For recommendation knowledge, we constructed subsets for user profiling (UProfile) and item title matching (ITMatching) based on the Amazon-Book (McAuley et al. 2015) to evaluate LLM’s comprehension of user and item textual description [^1] The evaluation details are provided in the Appendix.. As shown in Fig. 1, the LLM’s performance on all five datasets significantly decreased after fine-tuning on Amazon-Book using the prompt in TALLRec (Bao et al. 2023), indicating that its capacities for utilizing, reasoning, summarizing, profiling, and understanding recommendation texts have been compromised during fine-tuning.
Figure 2: Collaborative features interfering with LLM’s understanding of textual prompts. We use pre-trained Vicuna-7B as the ground truth. Our method avoids this interference.
(2) Existing methods endeavor to align collaborative information with the input space of LLMs by embedding collaborative features of users and items into textual prompts. Nevertheless, this practice can disrupt the LLM’s comprehension of the original text semantics. For instance, as shown in Fig. 2, while we expect Vicuna-7B without fine-tuning to repeat user and item description, the revised prompts fail to guide the LLM to generate the correct responses. Overall, existing studies limit general knowledge and collaborative information integration when adapting LLM to recommendation tasks.
To address this issue, we propose aligning collaborative features with the LLM’s parameter space by transforming them into plug-in weights to update LLM’s output. Unlike input space alignment, this parameter space alignment approach enables general LLM to directly leverage collaborative information as an adaption of each user-item pair for recommendation tasks, eliminating the need for fine-tuning or modifying the text prompts.
Specifically, the prediction for each user-item pair is divided into two parts: collaborative weights generation and text inference with collaborative weights. First, we utilize a pre-trained collaborative filtering model to extract user and item embeddings, which are then processed by a collaborative weight generator. The generator converts these embeddings to collaborative perceptual weights similar to “low-rank adaption” (LoRA (Hu et al. 2022)). We incorporate the collaborative weights into the LLM’s pre-trained weights to endow it with the capability to perceive collaborative information between the user and the item, allowing it to generate personalized predictions for each user-item pair with general text prompts. Importantly, our technique does not alter the LLM’s understanding of the original text, as it does not introduce new tokens. Instead, it updates LLM’s output with incremental weights. For tasks that do not involve collaborative information, our method seamlessly switches to utilizing the frozen backbone to infer textual prompts. Thereby, our procedure inherently addresses the aforementioned problems. To sum up, our contributions are summarized as follows:
  • We explore a novel paradigm for equipping LLMs with collaborative perception capability by merging collaborative weights into LLM’s pre-trained weights. It can be seen as Collaborative LoRA (CoRA), facilitating the cooperation of text and collaborative information. To the best of our knowledge, we are the first to endeavor collaborative perception in the parameter space of LLMs.
  • Following the CoRA method, we propose a collaborative weights generator that transforms collaborative features extracted from conventional recommenders into collaborative weights through a decoder-only architecture with a low-rank property.
  • Experimental results demonstrate the effectiveness of our approach, showing significant improvements over state-of-the-art LLMRec methods and traditional collaborative filtering methods on real-world datasets.
Related Work
Collaborative Filtering Models
Collaborative information is essential in the existing recommendation literature(Yuan et al. 2023; Liu et al. 2023b). In traditional personalized recommendation, collaborative filtering (CF) models are prevalent, leveraging collaborative information of users and items to generate predictions (Rendle et al. 2009; Guo, Zhang, and Yorke-Smith 2015). In these models, users and items are represented as latent factors fed into neural networks to model their interactions (He et al. 2020; Tang and Wang 2018; He et al. 2017). These studies achieved remarkable success in academia and industry, inspiring further exploration into collaborative information perception for LLMRec.
LLM for Recommendation
Given the impressive capabilities exhibited by LLMs, there is an increasing focus on exploring their potential applications in recommender systems. Techniques in this domain involve translating recommendation tasks into natural language tasks and adapting LLMs to generate recommendation results directly. These generative approaches can be divided into two paradigms based on whether parameters are tuned: non-tuning and tuning paradigms. The non-tuning paradigm assumes LLMs already possess the recommendation abilities and attempts to leverage their strong zero/few-shot abilities by introducing specific prompts (Liu et al. 2023a; Dai et al. 2023a; Mysore, McCallum, and Zamani 2023; Wang et al. 2023; Hou et al. 2023). In contrast, the tuning paradigm uses prompt learning or instruction tuning (Kang et al. 2023; Wang et al. 2022; Cui et al. 2022) to enhance LLM’s recommendation capacities.
A new trend is surfacing to include collaborative information in LLMs. Some studies focus on discovering user and item encoding methods to introduce new tokens through vocabulary expansion(Zheng et al. 2023; Hua et al. 2023; Rajput et al. 2023; Zheng et al. 2024). Others explore extracting collaborative information using latent factor models and aligning it with the input space of LLMs (Zhang et al. 2023, 2024b; Liao et al. 2023; Zhang and wei Wang 2023; Li et al. 2023d). While these methods exhibit strong performance, they constrain specific general capabilities of LLMs. Moreover, due to the hybrid prompting methods, collaborative features can disrupt original textual semantics. In this study, we propose CoRA, which generates low-rank collaborative weights like LoRA and injects them into LLM’s weights without extra inference overhead.
Preliminaries
In this section, we introduce the problem definition, the basic concepts, and the notations in this paper.
Problem Definition
We represent the interaction dataset as $\mathcal{D}=\{(u,i,y)|u\in\mathcal{U},i\in\mathcal{I}\}$ where $\mathcal{U}$ and $\mathcal{I}$ denote the set of users and items, respectively, with $y_{ui}=\{0,1\}$ indicating the interaction label. For an item $i$, there is a unique identifier and textual description $t_{i}$. For a user $u$, we construct the textual description from its historical interactions $t_{u}=\{t_{i}|i\in\mathcal{I}_{u}\}$. In this study, we aim to enable LLMs to leverage both collaborative embeddings $e_{u},e_{i}$ and textual descriptions $t_{u},t_{i}$ to predict whether a user $u$ will enjoy an item $i$ (i.e. $y_{ui}$).
Collaborative Filtering Models
To gather collaborative information, we are looking into utilizing CF methods, which represent users and items by latent factors (a.k.a. embeddings). User and item embeddings are variously calculated to capture collaborating relations precisely. The formulation of encoding a sample $(u,i,y)\in\mathcal{D}$ can be written as follows:
$\mathbf{e}_{u}=f_{\psi}(u;\mathcal{D});\ \mathbf{e}_{i}=f_{\psi}(i;\mathcal{D}),$ (1)
where $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d_{c}}$ denote the user’s and item’s representations with dimension $d_{c}$, $f_{\psi}(\cdot)$ denotes the encoding process, and $\psi$ represents model parameters. The user and item embeddings are then fed into the interaction prediction.
Large Language Model
Figure 3: Details of the LLM Decoder Block. (a) depicts the details of the LLM decoder block. (b) illustrates the multi-head self-attention module.
LLMs refer to a class of language models equipped with billions of parameters. As depicted in Fig. 3, their decoder block contains a multi-head self-attention module and a feed-forward network.
Self-attention contains four types of linear layers: query $W_{Q}$, key $W_{K}$, value $W_{V}$, and output $W_{O}\in\mathbb{R}^{h\times d}$, where $h$ and $d$ are dimensions of the hidden states and attention head. For each input sequence with $L$ tokens $X\in\mathbb{R}^{L\times h}$, it is multiplied by linear layers $W_{Q}$, $W_{K}$, $W_{V}$, obtaining $Q=XW_{Q}$, $K=XW_{K}$ and $V=XW_{V}$. Then, the self-attention operation is performed along the sequence dimension as follows:
$\mbox{Attention}(Q,K,V)=\mbox{softmax}(\frac{QK^{\top}}{\sqrt{d}})V.$ (2)
Then, the outputs parallel in different heads are concatenated and again projected by linear transformation matrix $W_{O}$, resulting in the final values.
Figure 4: Model architecture overview of our CoRA. The left path extracts user and item embeddings using a frozen collaborative model and converts them to collaborative weights through the generator. The right path fills the text fields in the prompt template, introducing textual descriptions for inference. Finally, the collaborative weights are merged into the LLM’s weights, enabling the LLM to perceive collaborative information without any fine-tuning or extra tokens in textual prompts.
Feed-forward Network is an MLP with two fully connected layers and one non-linear activation function. The formulation is as follows:
$\mbox{FFN}(x)=\sigma(xW_{1})W_{2},$ (3)
where $x$ is the input token, $W_{1}$ and $W_{2}$ are the weights of fully connected layers, and $\sigma$ is an activation function. Overall, there are five types of linear weights in a decoder block, including $W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$ in the self-attention module, and $W_{1}$, $W_{2}$ in the feed-forward module.
Methodology
In this section, we introduce our collaborative weights generator and CoRA method, which effectively integrates collaborative information into LLMs in the parameter space. The overall framework is shown in Fig. 4. First, we explain how to obtain collaborative weights from the pre-trained user and item embeddings through a collaborative weights generator. Then, we elaborate on how to equip LLM with collaborative perception capability with the help of generated collaborative weights, followed by a description of the prediction and training strategy.
Collaborative Weights Generator
Figure 5: Collaborative Weights Generator. Fig. (a) illustrates the pipeline of our collaborative weights generator. We set $k$ learnable perceptual queries, which interact with collaborative features through cross-attention modules to perceive collaborative information. Fig. (b) demonstrates that our approach is conceptually consistent with LoRA.
For a user-item pair $(u,i)$, we first obtain user and item embeddings $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d_{c}}$ from a pre-trained CF model and concatenate them as $[\mathbf{e}_{u},\mathbf{e}_{i}]$. To convert the features to collaborative weights $\Delta W\in\mathbb{R}^{h\times h}$, we propose a collaborative weights generator inspired by BLIP-2 (Li et al. 2023a).
Specifically, we design a decoder-only architecture to generate collaborative weights $\Delta W$. As shown in Fig. 5(a), the weights generator contains $N$ decoder blocks, each consisting of a self-attention module, a cross-attention module, and a feed-forward network. We set the hidden dimension of the generator to be equal to $d_{c}$. In the generator, $k$ learnable perceptual queries absorb collaborative information from different $k$ representation sub-spaces. In each decoder block, perceptual queries first interact with each other through self-attention modules, then interact with frozen collaborative features through cross-attention modules to perceive collaborative information, and finally go through a feed-forward network. Afterwards, we obtain $k$ collaboration-aware features $q_{c}\in\mathbb{R}^{d_{c}}$, which should be mapped and aligned with LLM’s parameter space. However, since $d_{c}\ll h\cdot h$, it will introduce a large number of parameters to directly map the dimension from $d_{c}$ to $h\cdot h$. Accordingly, a low-rank computation is considered in this process to reduce the computation costs. In particular, we adopt a shared linear layer $W_{share}\in\mathbb{R}^{d_{c}\times h\cdot r}$ to project all features $q_{c}\in\mathbb{R}^{d_{c}}$ to $W_{c}\in\mathbb{R}^{h\cdot r}$ as follows:
$W_{c}=q_{c}W_{share}$ (4)
where $W_{c}\in\mathbb{R}^{h\cdot r}$ is collaborative parameter and $r$ is the rank for collaborative weights.
Then, we reshape collaborative parameters $W_{c}$ as $h\times r$ and aggregate collaborative information from different sub-spaces with a mean-pooling operation. To obtain collaborative weights $\Delta W$, we ascend the dimension from $h\times r$ to $h\times h$ through a linear projection $W_{proj}\in\mathbb{R}^{r\times h}$ as follows:
$\Delta W=W_{c}W_{proj}.$ (5)
It is worth noting that the form of this low-rank calculation is equivalent to LoRA (Hu et al. 2022), where $W_{c}$ and $W_{proj}$ are the same as $\Delta W_{A}$ and $\Delta W_{B}$ in LoRA. Consequently, as illustrated in Fig. 5(b), our collaborative weights generator can be seen as a “Collaborative LoRA” generator.
Collaborative Perception in LLM
Existing LLMRec methods project user and item embeddings into the LLM’s input space to integrate collaborative and textual information. These approaches interfere with the LLM’s comprehension of the original textual prompts using its general capabilities, resulting in sub-optimal performance. To address this issue, we propose including collaborative information in LLM’s parameter space. Concisely, we convert collaborative information of input user-item pair $(u,i)$ to collaborative weights by the generator and merge them into LLM’s weights to enable LLM to perceive collaborative information without any fine-tuning or extra tokens in textual prompts. Without loss of generality, we utilize Vicuna-7B (Chiang et al. 2023) as the backbone LLM.
Concretely, for an input sample consisting of a user, an item, and their descriptions, we first adopt a CF model $f_{\psi}(\cdot)$ to extract the collaborative features $[\mathbf{e}_{u},\mathbf{e}_{i}]\in\mathbb{R}^{d_{c}}$. Then, we convert these collaborative features to collaborative weights $\Delta W\in\mathbb{R}^{h\times h}$ using the collaborative weights generator. To ensure that LLM can perceive collaborative information while preserving its general language capabilities, we sum $\Delta W$ with the frozen LLM’s weights directly as follows:
$\hat{W}=W+\Delta W=W+W_{c}W_{proj}$ (6)
where $W\in\mathbb{R}^{h\times h}$ denotes the overall weights of LLM.
By absorbing collaborative weights, the LLM inherently acquires the ability to perceive collaborative information. Additionally, this approach avoids interfering with or burdening the LLM’s text processing during inference. We can generate collaborative weights and incorporate them into any type of weights within LLM’s decoder block.
LLM prediction and Training Method
[…]
Table 1: Example of the used prompt template, using the same format as TALLRec (Bao et al. 2023).
Once LLM is equipped with collaborative perceptual capabilities, it can predict recommendations without additional fine-tuning. For an input sample $s=(u,i,t_{u},t_{i})$, the prediction generation can be formulated as:
$\hat{y}$ $=\mbox{LLM}(s)=\mbox{LLM}_{W+\Delta W}(p)$ (7)
$=\mbox{LLM}_{W+g_{\Theta}([f_{\psi}(u),f_{\psi}(i)])}(p)$
where $u,i$, and $t_{u},t_{i}$ represent the identifier and textual description of the user and item, respectively. $p$ is the prompt constructed with textual information $t_{u}$ and $t_{i}$ as shown in Tab. 1. $g_{\Theta}(\cdot)$ denotes the collaborative weights generator with parameters $\Theta$.
The only module that requires training is the collaborative weights generator, which converts pre-trained collaborative features to collaborative weights. We minimize prediction errors to optimize the parameters of the generator $\Theta$:
$\hat{\Theta}=argmin_{\Theta}\sum_{(u,i,y)\in\mathcal{D}}\ell(\hat{y},y),$ (8)
where $\ell(\cdot)$ denotes the loss function, which is implemented as the binary cross-entropy (BCE) loss.
Experiments
Experimental Settings
Dataset #Train #Valid #Test #User #Item
ML-1M 33,891 10,401 7,331 839 3,256
Amazon-Book 727,468 25,747 25,747 22,967 34,154
Table 2: Statistics of the processed datasets.
Dataset Amazon-Book ML-1M Method AUC UAUC Improve AUC UAUC Improve Collab. MF 0.7105 0.5543 14.04% 0.6486 0.6396 10.56% LightGCN 0.7026 0.5619 13.93% 0.5858 0.6512 15.68% SASRec 0.6675 0.5614 17.04% 0.7005 0.6734 3.65% LLMRec ICL 0.5180 0.5043 51.61% 0.5119 0.5178 38.37% Prompt4NR 0.6527 0.5011 25.10% 0.7027 0.6713 3.28% TALLRec 0.6583 0.4971 25.11% 0.7044 0.6741 3.31% LLMRec w/ Collab. PersonPrompt 0.7113 0.5596 13.44% 0.7014 0.6503 5.40% CoLLM-MF 0.8021 0.5782 5.14% 0.7028 0.6714 3.64% CoLLM-LGCN 0.7835 0.5663 7.48% 0.7164 0.6842 4.68% CoLLM-SAS 0.7538 0.5874 7.55% 0.7059 0.6531 4.84% BinLLM 0.8157 0.5724 4.83% 0.7132 0.6815 2.11% Ours CoRA-MF 0.8179 0.6262 - 0.7361 0.6884 - CoRA-LGCN 0.7886 0.5689 - 0.7128 0.6966 - CoRA-SAS 0.7677 0.5961 - 0.7019 0.6517 -
Table 3: Performance comparison of various models on Amazon-Book and ML-1M. ”Collab.” denotes collaborative recommendation methods. ”Improve” denotes the relative improvement of CoRA compared to baselines, averaged over the two metrics. All improvements are statistically significant, as determined by a paired t-test with $p\leq 0.05$.
Datasets. We adopt two widely-used public datasets for evaluation: ML-1M [^2] https://grouplens.org/datasets/movielens/1m/ and Amazon-Book [^3] https://jmcauley.ucsd.edu/data/amazon/index˙2014.html. For dataset processing, we adhere entirely to the setup of CoLLM (Zhang et al. 2023). The detailed statistics of the datasets are summarized in Tab. 2.
Baselines. To assess the effectiveness of CoRA, we compare it with three types of methods: conventional collaborative filtering methods (MF (Rendle et al. 2009), LightGCN (He et al. 2020), and SASRec (Kang and McAuley 2018)), LLMRec methods without collaborative information (ICL (Dai et al. 2023b), Prompt4NR (Zhang and wei Wang 2023), and TALLRec (Bao et al. 2023)), and LLMRec methods that consider collaborative information(PersonPrompt (Li, Zhang, and Chen 2022), CoLLM (Zhang et al. 2023), and BinLLM (Zhang et al. 2024b)).
Implementation Details. We use Vicuna-7B as the backbone LLM to implement our method. Binary Cross-Entropy (BCE) is employed as the objective function for optimizing all methods. AdamW (Loshchilov and Hutter 2017) optimizer is adopted for optimizing LLM, and Adam (Kingma and Ba 2014) optimizer for other methods. The detailed hyper-parameter settings and exploration can be found in the Appendix and the source code. We tune hyper-parameters according to the AUC metric on the validation dataset.
Regarding the collaborative weights generator, we set the hidden dimension and the embedding size of the collaborative model $d_{c}$ as 256 and the number of blocks $N$ as 8. The rank $r$ of perceptual weights is 16. The number of perceptual queries $k$ is 4. We insert collaborative weights $\Delta W$ on every decoder block of LLM. For better collaborative perceptual ability, we explore to insert $\Delta W$ for different types of weights in LLM, including $[W_{Q},W_{K},W_{V},W_{O},W_{FFN}]$. It is worth noting that the last linear layer $W_{proj}$ of the collaborative weights generator is zero-initialized for training stability as it is equivalent to the $\Delta W_{B}$ of LoRA weights.
For our CoRA, we implement it across three collaborative filtering methods to validate the effectiveness of our method in utilizing different types of collaborative information, denoted as CoRA-MF, CoRA-SASRec, and CoRA-LightGCN.
Evaluation Metrics. We employ two widely used metrics for click/rating prediction: AUC (Area under the ROC Curve, which measures the overall prediction accuracy) and UAUC (AUC averaged over users) (Liu et al. 2021) to evaluate the performance of our method and baselines.
(a) Amazon-Book Warm
(b) ML-1M Warm
(c) Amazon-Book Cold
(d) ML-1M Cold
Figure 6: Performance comparison in warm and cold scenarios on Amazon-Book and ML-1M. The left and right y-axis are AUC and UAUC, respectively.
Performance Comparison
Overall Performance. The performance is summarized in Tab. 3, from which we can find that our CoRA outperforms both collaborative filtering and LLMRec baselines, confirming the superiority of CoRA in leveraging collaborative information from traditional CF methods and general knowledge from LLM to make better predictions.
LLMRec with collaborative information surpasses traditional collaborative filtering methods, revealing that LLM’s extensive knowledge can be effectively generalized to recommendation systems. Furthermore, LLMRec methods that incorporate collaborative information show superior performance compared to those that do not, highlighting the significance of incorporating collaborative data into LLM for enhanced recommendations. Overall, integrating collaborative information with LLM’s general abilities represents a promising advancement in recommendation technology.
Our approach retains its advantages compared to LLMRec with collaborative information methods. When comparing the variants of CoRA and CoLLM based on different collaborative filtering models, CoRA consistently beats CoLLM. This indicates that aligning collaborative information with LLM’s parameter space is more effective in helping LLM perceive collaborative information and integrate it with linguistic competence.
Warm and Cold Performance. The main purpose of introducing collaborative information into LLMRec is to enhance the performance of LLMRec methods in warm-start scenarios. We divide the test data into warm and cold subsets based on the number of interactions. Without loss of generality, we compare the testing performance of MF, TALLRec, CoLLM, BinLLM, and CoRA on both warm and cold test subsets, as shown in Fig. 6.
In the warm scenario, TALLRec performs worse than MF because it does not consider collaborative information. Conversely, LLMRec methods that integrate collaborative information perform better, with CoRA achieving the best. This demonstrates that collaborative information is crucial for recommendations in warm-start scenarios, and our method is more effective in utilizing it.
In the cold start scenario, all LLMRec methods outperform MF, indicating that LLM’s universal capabilities can effectively alleviate the cold start problem by utilizing item textual information. Moreover, CoRA enhances the cold-start performance, suggesting that our method is superior to existing methods in integrating the collaborative knowledge of CF models and the general language knowledge of LLM.
Ablation Study
The effectiveness of integrating collaborative and textual information. To further validate the superiority of our method in enabling LLM to incorporate collaborative information, we construct variants of different LLMRec methods only with ID embeddings (i.e., removing textual descriptions of items). We treat TALLRec as a “Text-Only” variant of the three methods as a reference. The results are shown in Fig. 7, from which we can observe that our method outperforms the baselines when using only ID embeddings, confirming that our method can significantly enhance the collaborative perception ability of LLM. On this basis, the performance of all methods improved after involving item textual descriptions in most cases. Notably, the performance of CoLLM on ML-1M significantly declined after introducing textual information, indicating that aligning ID embeddings with LLM’s input space interferes with textual semantics and fails to combine their respective advantages. Our CoRA, on the other hand, shows a greater enhancement after incorporating text information, which we attribute to introducing collaborative information in LLM’s parameter space instead of the input space, fundamentally avoiding the interference problem.
(a) Amazon
(b) ML-1M
Figure 7: Performance of various variants. “ID-Only” refers to the removal of the item text. “w/ Text” represents adding item textual descriptions.
Weight Type Amazon-Book ML-1M AUC UAUC AUC UAUC qkvof 0.8141 0.6068 0.7312 0.6801 qkvo 0.8179 0.6262 0.7361 0.6884 qkv 0.7741 0.5747 0.6947 0.5933 qko 0.8091 0.5949 0.7111 0.5973 qk 0.7685 0.5644 0.6784 0.5887
Table 4: The impact of perceptual weighs type. q, k, v, and o denote query, key, value, and output linear weights in the self-attention module, respectively. f denotes the weights of feed-forward networks.
The type of equipped collaborative weights. As mentioned in Sec. Large Language Model, there are five types of weights in LLM’s decoder block, which are query, key, value, output, and feed-forward. We explore the impact of inserting collaborative weights for different types of LLM weights. As shown in Tab. 4, LLM equipped with collaborative weights for all types except feed-forward of weights performs best. Moreover, we observe that the performance of qkvo and qko is much better than qkv, suggesting that the output weights $W_{O}$ are essential for collaborative perception.
Conclusion
In this paper, we first explore the issues arising from aligning collaborative information in the input space of LLMs. To address them, we introduce CoRA, a novel paradigm that enables LLM to perceive collaborative information without fine-tuning or extra collaborative tokens. Our method converts collaborative information into LLM’s incremental weights through a collaborative weights generator, effectively integrating collaborative and textual information. Extensive experiments demonstrate the superiority of CoRA. For future work, we will expand experiments on other LLM backbones and recommendation tasks. Besides, we aim to extend our method to device-cloud collaborative learning.